Stock A has an expected return of 7%, a standard deviation of expected returns of 35%, a correlation coefficient with the market of -0.3, and a beta coefficient of -0.5. Stock B has an expected return of 12% a standard deviation of returns of 10%, a 0.7 correlation with the market, and a beta coefficient of 1.0. Which security is riskier? Why? (show your work)
Answer #1

The stock with the highest standard deviation to expected return ratio is the riskiest

Stock A standard deviation to expected return ratio = 0.35/0.07 = 5

Stock B standard deviation to expected return ratio = 0.10/0.12 = 0.833

Since, stock A standard deviation to expected return ratio is higher, Stock A is more risky